# CRNNet: Copy Recurrent Neural Network Structure Network

### v1

@@ 5 Experimental Result
We assessed the performance of various models, including Caml, RPGnet, CRNNet, and PLM-ICD. We evaluated these models using the top-50 label setting and trained the first three models with the same preprocessing method. Additionally, we incorporated a pretrained PLM-ICD (BERT) model in our analysis. The results presented in Table 1 lead to the following observations:
Among the compared models, CRNNet achieved the highest complication ratio. Notably, CRNNet’s complication ratio was 15.01% higher than PLM-ICD, the state-of-the-art method for EHR coding. This indicates that CRNNet effectively identifies complications associated with other codes. This may be attributed to CRNNet’s exploration of ICDs through a sequence of paths, as well as its utilization of information from previous codes to generate new ones.
@@
We conducted an ablation study to evaluate the effects of modules in CRNNet, shown in Table 2. Followings are ablation of CRNNet that we considered: (1) No copy represents CRNNet removed the copy module. We removed the copy module and used only the RNN structure model to generate codes. (2) No ARL represents CRNNet removed adversarial learning strategy. We removed the adversarial learning process to train our model. From the results, we obtained the following finding:
First, the ability of CRNNet to identify complication diseases decreased dramatically after removing the copy module. Specifically, the complication ratio in the result of CRNNet(no-copy) dropped by 5.44% respectively. The result shows that copy module plays a significant role in our method to identify complicating diseases. The copy module extracts information related to complicating diseases based on the former ICD in the path, which helps to generate complicating diseases.
Second, the ability of CRNNet to identify complicating diseases is still working better than PLM-ICD when copy module is not used. The complication ratio of PLM-ICD is 42.29% and that of CRNNet(no-copy) is 51.86%, which indicates that the RNN structure is also helpful in identifying complicating diseases, as the ICD on timestamp $t+1$ is generated based on the ICD on timestamp $t$.
Finally, the adversarial reinforcement learning (ARL) does not significantly improve the performance of CRNNet, but it slightly improves its ability to identify complicating diseases. This is due to adversarial reinforcement learning helps CRNNet generate codes paths similar to the ground truth paths.
@@ 6 CONCLUSION AND FUTURE WORK
In this paper, we propose an innovative approach for Electronic Health Record (EHR) coding, specifically targeting the identification of complicating diseases. We frame this as a path generation task and introduce the Copy Recurrent Neural Network Structure Network (CRNNet). CRNNet consists of a EHR-ICD fusion module designed to fuse the representation of EHRs and code path. An adversarial learning framework is used for training stage. Our experimental results on the MIMIC-III dataset reveal that CRNNet surpasses existing methods in detecting complicating diseases, and its performance excels when using the CNN encoder method. The performance improvement is attributed to the implementation of the copy module and the adversarial reinforcement learning mechanisms.
Nonetheless, CRNNet’s limitations lie in its improvement of recall and AUC. While it performs well in detecting complications, it does not focus on enhancing basic performance metrics. Additionally, although it is based on a CNN encoder, transformer methods are now considered a more efficient way to encode EHRs. In future research, we will focus on integrating transformer methods and improving performance with the aid of complication detection, which is a crucial task in the EHR coding area.
7 DATA AVAILABILITY
We make our code and dataset publicly available at:
https://anonymous.4open.science/r/CRNNNet-D780

### v2

@@ 5 Experimental Result
We assessed the performance of various models, including Caml, RPGnet, ComplicaCode, and PLM-ICD. We evaluated these models using the top-50 label setting and trained the first three models with the same preprocessing method. Additionally, we incorporated a pretrained PLM-ICD (BERT) model in our analysis. The results presented in Table 1 lead to the following observations:
(1)Among the compared models, ComplicaCode achieved the highest complication ratio. Notably, ComplicaCode’s complication ratio was 15.01% higher than PLM-ICD, the state-of-the-art method for EHR coding. This indicates that ComplicaCode effectively identifies complications associated with other codes. This may be attributed to ComplicaCode’s exploration of ICDs through a sequence of paths, as well as its utilization of information from previous codes to generate new ones.
(2)PLM-ICD is powered by BERT, a significantly more advanced approach for extracting text representations, utilizing a pretrained model that benefits from training on an larger dataset. This superiority underscores the correlation between a high-quality model and its capability in detecting complex diseases. Moreover, the rate of complication detection serves as a crucial metric for assessing a model’s performance. While our model may lag behind PLM-ICD in terms of Jaccard, it excels in complication detection, a testament to both the efficacy of our model and the promise of our proposed module.
@@
We conducted an ablation study to evaluate the effects of modules in ComplicaCode, shown in Table 2. Followings are ablation of ComplicaCode that we considered: (1) No copy represents ComplicaCode removed the copy module. We removed the copy module and used only the RNN structure model to generate codes. (2) No ARL represents ComplicaCode removed adversarial learning strategy. We removed the adversarial learning process to train our model. From the results, we obtained the following finding:
First, the ability of ComplicaCode to identify complication diseases decreased dramatically after removing the copy module. Specifically, the complication ratio in the result of ComplicaCode(no-copy) dropped by 5.44% respectively. The result shows that copy module plays a significant role in our method to identify complicating diseases. The copy module extracts information related to complicating diseases based on the former ICD in the path, which helps to generate complicating diseases.
Second, the ability of ComplicaCode to identify complicating diseases is still working better than PLM-ICD when copy module is not used. The complication ratio of PLM-ICD is 42.29% and that of ComplicaCode(no-copy) is 51.86%, which indicates that the RNN structure is also helpful in identifying complicating diseases, as the ICD on timestamp $t+1$ is generated based on the ICD on timestamp $t$.
Finally, the adversarial reinforcement learning (ARL) does not significantly improve the performance of ComplicaCode, but it slightly improves its ability to identify complicating diseases. This is due to adversarial reinforcement learning helps ComplicaCode generate codes paths similar to the ground truth paths.
@@ 6 CONCLUSION AND FUTURE WORK
In this paper, we propose an innovative approach for Electronic Health Record (EHR) coding, specifically targeting the identification of complicating diseases. We frame this as a path generation task and introduce the ComplicaCode. ComplicaCode consists of a EHR-ICD fusion module designed to fuse the representation of EHRs and code path. An adversarial learning framework is used for training stage. Our experimental results on the MIMIC-III dataset reveal that ComplicaCode surpasses existing methods in detecting complicating diseases, and its performance excels when using the CNN encoder method. The performance improvement is attributed to the implementation of the copy module and the adversarial reinforcement learning mechanisms.